OpenStudy (anonymous):

A rectangular swimming pool is 50 feet long and 25 feet wide. What is the area of the bottom of the swimming pool? ft2

OpenStudy (stormer15):

50 * 25 = ?

OpenStudy (anonymous):

1250

OpenStudy (stormer15):

so that's the answer 1250 ft^2

OpenStudy (anonymous):

A rectangular picture is 7 inches long and 5 inches wide. What is the area of the picture? A. 35 in² B. 35 in. C. 24 in. D. 24 in²

OpenStudy (stormer15):

7 * 5 = ?

OpenStudy (anonymous):

35

OpenStudy (stormer15):

and remember the answer for area has unit^2

OpenStudy (anonymous):

What is the area of this rectangular television screen? A. 65 in. B. 130 in² C. 966 in² D. 966 in.

OpenStudy (stormer15):

did I give direct answer? LOL i told them what they were supposed to multiply

OpenStudy (stormer15):

42 * 23 = ?